### v1

@@ 3.3 Small world network
We note that even though the use of tensor product algorithms allows us to compute likelihoods (4) faster and more accurately, Bayesian inference of a contact network in a fully black–box setting remains a challenging task, as we see from experiments in Sec. 3.1 and Sec. 3.2.
In this section we present a preliminary experiment where we assume some prior knowledge of the contact network, which allows us to reduce the number of unknown parameters even for a larger number of network nodes. Specifically, we assume that the contact network is from a family of small–world networks [74], which is shown in Figure 6(a). It consists of nodes which are arranged as a loop and connected with a double bond, where each node is connected to nodes and where we assume that indices go around the circle, so and when necessary. The main loop is rewired, i.e a certain link is removed and replaced with a link to a random node which provides additional connectivity. For this experiment we assumed that the ground truth network contains a single rewired link i.e. the link is removed and replaced with We then proceed to infer this network, assuming that we know it is from the set of small–world networks with a single rewired link which we denote The problem therefore reduces to finding only two parameters, and and the search space shrinks from to only possible grids.
Inferring a network from a known class can be formulated as Bayesian optimisation (4) on a class of networks parametrised by small number of parameters. This removes our main computational challenge related to high dimensionality of the search space and allows us to solve this problem directly. We generated data samples by simulating the –SIS epidemic on a ground truth contact network using parameters and for the duration of time units, and re-sampled the data to a uniform grid with the time step hence creating records for each data sample. Using tAMEn algorithm to model the evolution of epidemic on –node networks, we were able to compute the likelihoods for all grids We then computed the average contrast for all as shown in Fig. 6(b). The results show that for all which ensures that the ground truth network is a unique global maximum of the Bayesian optimisation problem (4).
@@ 4 Discussion
As the optimisation gets closer to the likelihoods increase and the presence of steep local maxima slows down the convergence towards the global one. In this area high contrast ratios are undesirable as they make it harder for the MCMC algorithm to get out of local maxima. This can be addressed by using only a part of the available data to compute likelihoods (4) and/or tempering of to simplify the high–dimensional landscape for the optimisation. This has been used successfully for sampling from concentrated distributions of continuous random variables [14].
@@
We also explored the potential of tensor product algorithms for tackling the network likelihood optimisation. However, these attempts so far were less efficient than the MCMC algorithm (in particular the MCMC without replacement). The TT-Cross algorithm [61] and its extensions [20] are used to compute a TT approximation to a black-box tensor by drawing a few adaptive samples from it using the maximum volume algorithm [32] or a greedy version thereof [20]. These maximum volume samples are expected to be good candidates for the maximum absolute value of the tensor [32, 68]. However, the maximum volume algorithm requires all elements of a TT core, which must be drawn as full columns from the tensor, including elements which are known to be far from the maximum. MCMC probes only one element at a time, and can skip such unnecessary calculations. In numerical experiments with the linear chain, MCMC was systematically faster and more accurate compared to the TT-Cross maximizer, albeit by a modest margin (1-2 contacts). Tempering the likelihood to reduce its TT ranks and caching its values (which are often repeated in the TT-Cross) may make this approach faster in terms of the actual CPU time.
@@
Another tensor optimiser proposed recently is PROTES [7], a probabilistic method similar to genetic algorithms. In each iteration, this algorithm draws candidate optima as random samples from a probability distribution function in the TT format, which is in turn updated by a stochastic gradient ascent maximizing the probability of drawing samples with the largest values of the sought function out of the candidates. The default parameters proposed in [7] are and . Compared to our budget of function evaluations, this corresponds to only stochastic gradient ascent iterations, which are clearly insufficient and produce an almost random network. Taking in the order of (and hence below ) is uncompetitive too, since a few tens of iterations cannot compensate for a more random stochastic gradient due to a smaller . However, it may be reasonable to use such an algorithm to fine-tune a previous TT approximation of the likelihood to new data.

### v2

@@ 3.4 Small world network
We note that even though the use of tensor product algorithms allows us to compute likelihoods (4) faster and more accurately, the exact Bayesian inference of a contact network in a fully black–box setting remains a challenging task, as we see from experiments in Section 3.2 and Section 3.3.
In this section we present a preliminary experiment where we assume some prior knowledge of the contact network, which allows us to reduce the number of unknown parameters even for a larger number of network nodes. Specifically, we assume that the contact network is from a family of small–world networks [76], which is shown in Fig. 8(a). It consists of nodes which are arranged as a loop and connected with a double bond, where each node is connected to nodes and where we assume that indices go around the circle, so and when necessary. The main loop is rewired, i.e a certain link is removed and replaced with a link to a random node which provides additional connectivity. For this experiment we assumed that the ground truth network contains a single rewired link i.e. the link is removed and replaced with We then proceed to infer this network, assuming that we know it is from the set of small–world networks with a single rewired link which we denote The problem therefore reduces to finding only two parameters, and and the search space shrinks from to only possible grids.
Inferring a network from a known class can be formulated as Bayesian optimisation (4) on a class of networks parameterised by a small number of parameters. This removes our main computational challenge related to high dimensionality of the search space and allows us to solve this problem directly. We generated data samples by simulating the –SIS epidemic on a ground truth contact network using parameters and for the duration of time units, and re-sampled the data to a uniform grid with the time step hence creating records for each data sample. Using tAMEn algorithm to model the evolution of epidemic on –node networks, we were able to compute the likelihoods for all grids We then computed the average contrast for all as shown in Fig. 8(b). The results show that for all which ensures that the ground truth network is a unique global maximum of the Bayesian optimisation problem (4).
@@ 4 Discussion
As the optimisation gets closer to the likelihoods increase and the presence of steep local maxima slows down the convergence towards the global one. In this area high contrast ratios are undesirable as they make it harder for the MCMC algorithm to escape local maxima. By preliminary experiments demonstrated in this paper we show that this can be addressed tempering of to simplify the high–dimensional landscape for the optimisation. Another idea is to use only a part of the available data to compute likelihoods (4), which has been used successfully for sampling from concentrated distributions of continuous random variables [77].
@@
We also explored the potential of tensor product algorithms for tackling the network likelihood optimisation. However, these attempts so far were less efficient than the MCMC algorithm (in particular the MCMC without replacement). The TT-Cross algorithm [78] and its extensions [79] are used to compute a TT approximation to a black-box tensor by drawing a few adaptive samples from it using the maximum volume algorithm [80] or a greedy version thereof [79]. These maximum volume samples are expected to be good candidates for the maximum absolute value of the tensor [80, 81]. However, the maximum volume algorithm requires all elements of a TT core, which must be drawn as full columns from the tensor, including elements which are known to be far from the maximum. MCMC probes only one element at a time, and can skip such unnecessary calculations. In numerical experiments with the linear chain, MCMC was systematically faster and more accurate compared to the TT-Cross maximiser, albeit by a modest margin (1-2 contacts). Tempering the likelihood to reduce its TT ranks and caching its values (which are often repeated in the TT-Cross) may make this approach faster in terms of the actual CPU time.
@@
Another tensor optimiser proposed recently is PROTES [82], a probabilistic method similar to genetic algorithms. In each iteration, this algorithm draws candidate optima as random samples from a probability distribution function in the TT format, which is in turn updated by a stochastic gradient ascent maximising the probability of drawing samples with the largest values of the sought function out of the candidates. The default parameters proposed in [82] are and . Compared to our budget of function evaluations, this corresponds to only stochastic gradient ascent iterations, which are clearly insufficient and produce an almost random network. Taking in the order of (and hence ) is uncompetitive too, since a few tens of iterations cannot compensate for a more random stochastic gradient due to a smaller . However, it may be reasonable to use such an algorithm to fine-tune a previous TT approximation of the likelihood to new data.
Choosing a more informative prior on the network may aid the inference. We have already stepped away from a fully uniform prior in the small world example, where we sought only one rewiring instead of the state of all links. Penalising improbable or redundant links with a low prior probability may be beneficial for more general networks as well.
Potentially, it may be possible to use all MCMC points to compute posterior expectations rather than the MLE/MAP. However, this would be difficult for network identification for two reasons. First, accurate sampling would need much more likelihood evaluations (and hence CPU time) to decorrelate the Markov chain, whereas the MLE/MAP can be found in a few thousand samples. Secondly, the expected state would be a real-valued instead of binary vector, and require ambiguous post-processing to convert it into a network. Mitigation of these obstacles can be a matter of a future research.
Abbreviations
SIR = susceptible–infected–recovered epidemic model, see [1]
SIS = susceptible–infected–susceptible epidemic model, see e.g. [83]
ODE = ordinary differential equation
CME = chemical master equation, see [7]
MLE = maximum likelihood estimate, see e.g. [84]
MAP = maximum a posteriori estimate, see e.g. [84]
SSA = stochastic simulation algorithm, see [8]
MCMC = Markov chain Monte Carlo [74]
MCMC-noR = a version of MCMC without replacement, see Sec. 2.6
TT = tensor train format, see [64]
MPS = matrix product states, see e.g. [73]
DMRG = density matrix renormalisation group algorithm, see [85]
AMEn = alternating minimal energy method, see [55]
tAMEn = time–dependent AMEn, see [56]
CPU time = central processing unit time, also known as the wallclock time
Declarations
Funding Sergey Dolgov was supported by the Engineering and Physical Sciences Research Council (EPSRC) New Investigator Award EP/T031255/1. Dmitry Savostyanov was supported by the Leverhulme Trust Research Fellowship RF-2021-258 at the initial stage of this work. Dmitry Savostyanov would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Discretization and recovery in high-dimensional spaces, where work on the revised version of this paper was undertaken. This work was partly supported by EPSRC grant no EP/R014604/1 and by a grant from the Simons Foundation. Funders took no part in study design; in the collection, analysis and interpretation of data; in the writing of the report; and in the decision to submit the article for publication.
Conflict of interest Authors declare no competing interests exist.
Ethics approval and consent to participate Not applicable.
Consent for publication Not applicable.
Data and code availability Numerical experiments in this work are based on synthetic randomly generated datasets. All data and code required to reproduce experiments are available on github.com/savostyanov/ttsir.
Materials availability Not applicable.
Authors’ contribution Sergey Dolgov developed software, performed numerical experiments, analysed the results, and contributed to writing the manuscript. Dmitry Savostyanov designed the work, analysed the results, designed visualisations, and was a major contributor to writing the manuscript. Both authors read and approved the final manuscript.